OpenStudy (anonymous):

If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its center point is help plez

OpenStudy (freckles):

(x-h)^2+(y-k)^2=r^2 (h,k) is center you have (x-(-5))^2+(y-7)^2=36 identify h and k